OpenStudy (anonymous):

question 1 A theater has 500 seats. three-fourth of seats are filled how many seats are filled? show steps how you got the answer 1/2 / 1/2 = show steps how you got that answer @zepdrix

OpenStudy (anonymous):

@zaynab123 @ZakkDanger @popcornfun

OpenStudy (anonymous):

question 2. 1/2 / 1/2 show stps how you got the answer

zepdrix (zepdrix):

Take the 500 seats. Break it up into 4 equal groups. Then add up 3 of those groups.

OpenStudy (anonymous):

375

zepdrix (zepdrix):

125 in each group? 3 of those add up to 375? Ok good job :)

zepdrix (zepdrix):

The second question is a little hard to read. Is it formatted like this?\[\large\rm \frac{\frac{1}{2}}{\frac{1}{2}}\]

OpenStudy (anonymous):

well it's 1/2 divided by 1/2 and im supposed to find the answer and explain how i got it

zepdrix (zepdrix):

Bah I did that backwards hehe, one sec

zepdrix (zepdrix):

If you had been given this question:\[\large\rm 2\div\frac{1}{2}\]It's essentially asking you, how many 1/2's can fit into the number 2. It ends up being 4, yes? Because 1/2 + 1/2 + 1/2 + 1/2 = 2.

zepdrix (zepdrix):

So for your problem, it's simply this:\[\large\rm \frac{1}{2}\div\frac{1}{2}\]How many 1/2's can fit into 1/2?

OpenStudy (anonymous):

1

zepdrix (zepdrix):

Wow the page just refreshed :( Erased everything I was writing.. ugh Anyway, yes. 1 is your answer :)

OpenStudy (anonymous):

but how do you get the answer @zepdrix

zepdrix (zepdrix):

We can apply one of two different rules. The first one is pretty simple: Whenever you divide a non-zero quantity by itself, the result is always 1.\[\large\rm 2\div 2=1\]\[\large\rm 5\div5=1\]\[\large\rm 37.23\div37.23=1\]\[\large\rm \frac{1}{8}\div\frac{1}{8}=1\]

zepdrix (zepdrix):

The other explanation you could use, I'm not sure if you've learned this yet though...

OpenStudy (anonymous):

thankyou soo much ur a big help 1 medal for you

zepdrix (zepdrix):

We have this "Keep Change Flip" rule for dealing with division of fractions. Example:\[\large\rm \frac{2}{3}\div\frac{5}{8}\quad=\frac{2}{3}\times\frac{8}{5}\]We `Keep` the top fraction the same. We `Change` the operation from division to multiplication. We `Flip` the bottom fraction. And so in your problem, it would have become,\[\large\rm \frac{1}{2}\div\frac{1}{2}\quad=\frac{1}{2}\times\frac{2}{1}\quad=\frac{2}{2}\quad=1\]

zepdrix (zepdrix):

That's a little bit of a fancier way to do it though :)